OpenStudy (anonymous):

Let c be any constant number. Which of the following will always be perpendicular to −4x + y = −2? y equals four x plus c y equals negative four x plus c y equals negative one over four times x plus c y equals one over four times x plus c

OpenStudy (lalaly):

the third one

myininaya (myininaya):

the 3rd one is correct

myininaya (myininaya):

-4x+y=-2 y=4x-2 perpendicular line has form y=-x/4-C

OpenStudy (anonymous):

it's the third one XD

OpenStudy (aroub):

perpendicular= reciprocal of the slope u have.. lol! yup its totally the third one!